OpenStudy (anonymous):

Multiply x to the 1 third power times x to the 3 sevenths power

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large x^{\frac{1}{3} }\times x^{\frac{3}{7} } \)